OpenStudy (anonymous):

Factor: 5x2 + 4x - 1

OpenStudy (wasiqss):

5x^2+5x-x-1 5x(x+1)-1(x+1) (5x-1)(x+1)

OpenStudy (anonymous):

\[5x ^{2}-x+5x-1=0,5x ^{2}+5x-x-1=0,5x(x+1)-1(x+1)=0,(5x-1)(x+1)=0\]